OpenStudy (laddiusmaximus):

f(x)= sin^2x i get using the product rule, but I get stuck after 2sinx cosx

OpenStudy (anonymous):

what are you trying to do?

OpenStudy (mimi_x3):

\[f(x) = (sinx)^2\] \[f'(x) = 2(sinx)*(cosx)\] So the trig identity is: \[\sin2x= 2sinxcosx\]

OpenStudy (anonymous):

also (sinx)^2 = 1/2*(1-cos2x)

OpenStudy (mimi_x3):

Yeah, but that is hard to differentiate i guess..

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

you don't have to use product rule

OpenStudy (anonymous):

just the chain for cos(2x)

OpenStudy (mimi_x3):

I didn't use the product rule for that, i think i used the chain rule..

OpenStudy (laddiusmaximus):

I understand that. I know what the chain rule is he hasnt taught it to us yet. the example demonstrated the product rule.

OpenStudy (anonymous):

remember the formula above for when you're going to integrate sin^2(x)

OpenStudy (mimi_x3):

@math: um, this is differentiate, i think you got kind of confused

OpenStudy (anonymous):

not really

OpenStudy (laddiusmaximus):

how would i solve this using only the product rule?

OpenStudy (anonymous):

(sinx)^2 = 1-cosx*cosx

OpenStudy (anonymous):

now you can use the product rule